the sum of leah's and brandon's ages is 34. five years ago leah was 3 times as old as brandon was. how old are leah and brandon now?
Answer by stanbon(75887)

the sum of leah's and brandon's ages is 34. five years ago leah was 3 times as old as brandon was. how old are leah and brandon now?
-----
Equations:
L + B = 34
(L-5) = 3(B-5)
-----
L = 34-B
Substitute for "L" and solve for "B":
(34-B-5) = 3B-15
29-B = 3B-15
4B = 44
B = 11 yrs (Brandon's age now)
----
L = 34-B = 23 yrs (Leah's age now)
=========================
Cheers,
Stan H.